Using imlicit differentiation, determine the derivative of "cos(x-2y) = sin(2x+y)"


1
Expert's answer
2021-06-08T14:48:22-0400
"\cos(x-2y)=\sin(2x+y)"

Differentiate both sides with respect to "x"


"\dfrac{d}{dx}(\cos(x-2y))=\dfrac{d}{dx}(\sin(2x+y))"

Use the Chain Rule


"-\sin(x-2y)(1-2\dfrac{dy}{dx})=\cos(2x+y)(2+\dfrac{dy}{dx})"

Solve for "\dfrac{dy}{dx}"


"\dfrac{dy}{dx}=\dfrac{\sin(x-2y)+2\cos(2x+y)}{2\sin(x-2y)-\cos(2x+y)}"